Run Planning Discussion 3 July, 2015

Magnet training. Will take beam with pion beams to fill out remainder of alignment requirements.

Daily schedule

July 3
estimated time 10pm - 6am
MOM R. Bayes
Lead shifter Mariyan Bogomilov
Shifter 2 Dan Kaplan
BLOC A. Dobbs
SOC C. Pidcott
Physics Shifter F. Drielsma
Aim Magnet commissioning
Beam? No
Additional Personnel

Hall Status

  • The priority is placed on the magnet preparation activities.
  • Cooling SSD and SSD is "complete".
  • TOF1 will be moved out of the beam line for required Tracker repairs Friday morning. The leak was found too late to put TOF1 in. Data taken using a TOF0 trigger would be of questionable value. Thus we will not take data tonight.

Shift plan:

  • 10pm
    Shifters arrive at hall prepared to monitor the solenoid cold mass pressures anticipating further instability in the instrumentation.

Data Requested

  • The total estimated data requirement for the tracker to tracker alignment was 300000 tracks.
  • The total number of TOF2 events collected since 20 June (when the tracker was stabilized so that useful data could be collected) is 877156. This number is based on the count from the CDB including both muon and pion runs. This number omits runs with no end of run transmission (including run 7168 which was composed of 70k TOF2 triggers generated from a muon beam).
  • This suggests that any beam we take tonight will not provide a significant increase in the generated statistics.

Beamline Configuration

  • No beams will be generated tonight.

Notes for shifters

  • Please continue to log information into the e-log independent of run information.

Phone details for 6pm daily meeting

Phone access : ReadyTalk number

US 866-740-1260

UK 0800-496-0576

Swiss 0800-70-53-51

Italy 800182592

Access Code: 2155402

Updated by Bayes, Ryan almost 9 years ago ยท 8 revisions